# Large Language Models and Games: A Survey and Roadmap

### v1

@@ 3.3 Player Assistant
A somewhat less explored role for LLMs in games is that of a player assistant: an interactive agent intended to enrich or guide the player experience in some way. This could be in the form of a sequence of tutorial-style tips, a character that does not causally interact with the game world at all, or an agent that is able to interact with the game world at a similar level as the player. Existing games make use of player assistants in different ways. For example, in The Sims (Electronic Arts, 2000) a disembodied assistant provides tips specific to the game context via dialogue boxes. Civilization VI (Firaxis Games, 2016) uses different assistants with a visual depiction to provide similar advice suggesting the best build option according to their idiosyncratic heuristic; they may thus alleviate some decision-making from the player. In management games, AI may automate menial tasks such as assigning jobs to a planet’s population in Stellaris (Paradox Interactive, 2016); this assistance reduces cognitive load from the player, but the player can always micro-manage this task if they wish.
LLMs will likely be appealing as player assistants given their potential expressive and conversational capacities. Coupled with either LLM-powered or heuristic-based methods for finding the best policy or action given the current game context (see Section 3.1), an LLM-powered player assistant can construct the suggested action and the explanation as a natural language utterance delivered by a disembodied or embodied agent, expressed with a corresponding sentiment, and manifested through its body stance, gestures and facial expression. Similarly, LLMs may assist the player by undertaking some minor tasks in the game via a tailored smaller role as “player” within that smaller task description (see Section 3.1). The potential of LLM-powered player assistants is not explored in current research, to the best of our knowledge. We highlight the potential of this future application in Section 4.
@@ 3.4 Commentator/Reteller
The vision of automated “let’s play”-style commentary generation is not new. It was proposed in [39] and implemented via classical machine learning methods, to limited success. In [40], an LSTM with text, vision and game-state input was trained to generate characters for a commentary script in a racing game. Results of this approach featured repetitive and context-irrelevant generated text. LSTMs were also used by [41] to generate text, at a character level, for Getting Over It With Bennett Foddy (Bennett Foddy, 2017), a challenging side-scrolling climbing game.
LLMs for commentary are also explored by Renella and Eger [42], who argue that LLMs could assist games streamers (e.g. on Twitch) while the streamer multitasks gameplay with audience interaction. The authors develop a pipeline for automatically commenting upon League of Legends (Riot Games, 2009) games. They take a multi-phased approach, training a model on hand-annotated data to recognize key events, then prompting ChatGPT to generate zero-shot commentary on these events in the style of a particular (known) fictional character, and finally sending the generated text through the FakeYou [^2] https://fakeyou.com/ API to be voiced in the timbre of this same character. For example, once the event detection model has identified an enemy double kill in a particular frame, ChatGPT responds in the style of Rick Sanchez, from Rick and Morty (Cartoon Network, 2013): “What the heck?! That enemy team just got a double kill! I can’t believe it! They must be pretty good! I better watch out for them!” An additional loop buffers detected events—for example, delaying commentary on a double kill in case it should escalate into a triple kill, or prioritizing among a quick barrage of events—and prompts ChatGPT to generate random fillers, such as thanking (fictional) new subscribers.
@@ 3.5 Game Master
A Game Master (GM) in table-top role-playing games (TTRPGs) is the person who creates the plot of a game, its characters, and narrative. GMs wear many hats during the course of the game session [43], but also before and after the session by preparing and adapting adventures and stories and checking in with the group [44]. Digital games have mostly prescripted stories or level progressions, and their players have a restricted range of affordances, compared to TTRPG players whose actions are only limited by their imagination. Similarly, the story told around the table can take any direction. Since human GMs mostly communicate about the gameworld, story, game state and action resolutions via natural language (although props such as maps, miniatures, hand-outs are also common), the potential of LLMs as a GM is often mentioned both in research circles and TTRPG discussion boards [45]. LLMs as GMs also open the potential for solo play, while a TTRPG requires at least one player and the human GM.
@@
One of the first notable text-based adventures managed by a fine-tuned version of GPT-2 is AI Dungeon [46]. AI Dungeon [^3] https://play.aidungeon.com/ is an online interactive chat-based storytelling application where the player takes actions through semantic input alone. The LLM continues the story based on the player’s input, in the fashion of a human GM. The game has evolved since its creation to make use of more recent LLM models, which the player can choose from before starting a play session. Different game world settings are also offered, and players are also able to share the stories they create. Similar games have emerged online ever since [^4] https://koboldai.net, https://www.hiddendoor.co, and a freely available code repository, Kobold AI Client [^5] https://github.com/KoboldAI/KoboldAI-Client, allows a local or remote installation of a client for such LLM-run games. Some of these games also use Stable Diffusion text-to-image models [47] to generate visuals accompanying different parts of the narrative.
@@
In lieu of replacing a human GM, LLMs have also been employed as GM assistants. CALYPSO [48] is a set of tools running on a Discord server which the GM can query either to generate random encounters, brainstorm ideas, or alternatively chat with a fictional character in a Dungeons & Dragons (TSR, 1974) TTRPG setting. CALYPSO highlights that hallucinations of GPT-3 can have both positive effects when it generates plausible details not included in descriptions published in the original game manual (e.g. the shapes of creatures’ eyes) and also negative effects when the created details are outright incorrect (e.g. describing the wings of a canonically wingless creature). In addition, the model’s preconditioning to avoid racial bias was found to occasionally prevent it from generating racial details of fantasy creatures in the game. Other work used smaller GPT models to improvise in-game conversations [49] by monitoring and transcribing verbal exchanges between the GM and the players, and attempting to generate appropriate responses. This example was integrated into “Shoelace” [50], which is itself a GM assisting tool helping with content lookup by creating a node-based plan of the game narrative and encounters. The versatility of LLMs given their ability to rapidly process text input paves the way for their integration into the multitude of existing tools and aids for human GMs.
@@ 3.6 Game Mechanic
Games can also be built around a specific mechanic that relies on LLMs, similar to the AI-based game design patterns identified in [52]. An obvious mechanic revolves around the social interactions facilitated by LLM-powered conversational NPCs. In this vein, [53] has employed LLMs to populate a virtual village with 25 characters, enabling them to communicate and engage in social behavior within a sandbox environment. Players were able to interact with these agents using a text-based interface. The environment state and actions of each agent were stored in a language-based format and summarized in order to retain knowledge for each agent when prompting for its actions. This led to emerging believable social interactions, such as the agents spontaneously inviting other agents to a party which one of them was organizing. Similarly, GoodAI are developing the AI people video game which operates as a sandbox simulation where LLM-powered NPCs “interact with each other and their environment, forming relationships and displaying emotions” [28]. The player can interact with the agents via natural language chat, triggering reactions and potentially disrupting the relationship between NPCs (see Fig. 1).
@@
Another strength of LLMs is language synthesis, which is leveraged by Infinite Craft [^7] https://neal.fun/infinite-craft/. Infinite Craft is an “alchemy” game, in which the player combines elements to produce new ones (see Fig. 3). In Infinite Craft, the player begins with a set of core elements (water, fire, wind and earth). But while the former have a set of interactions defined manually by the designer, Infinite Craft prompts LLaMA 2 [12] to imagine the product of the combination of these elements [55]. Judging from gameplay, it appears that for each distinct combination, LLaMA is prompted to produce the result only once, with the product stored in a database for future reference. Thus seemingly anything in the language model’s vocabulary might “emerge” from the combination of these elements, including all 50 states [^8] https://x.com/FeralFlex/status/1758332430136615298, “Dream” [^9] https://x.com/slutzsmp/status/1760394123243135169, and the fictional “Super Stonedosaurus Tacosaurus Rex” [^10] https://x.com/pvtspicy/status/1759316982984237139. On occasion, the model can choose to return one of the combined elements, or refuse to combine (in particular very lengthy or complex) elements.
@@ 3.7 Automated Designer
Prior work in PCG has demonstrated that tile-based game levels can be reliably generated with sequence-based prediction models (e.g. LSTMs) from a modest set of examples, by treating such levels as linear sequences of tile types in raster order [57]. More recently, modern LLMs have been leveraged in a similar way to generally greater success. In [58], a GPT-2 model was fine-tuned on a large dataset of Sokoban (Thinking Rabbit, 1982) levels and, at test time, samples from the model to produce novel puzzles (see Figure 4). Interestingly, their results indicate that while the GPT-2 model struggles when the fine-tuning dataset is restricted in size, GPT-3 (and, presumably, larger models released since then) are better able to accommodate limited training sets.
@@
A similar approach, MarioGPT, trains a GPT-2 model on a relatively small dataset of Super Mario Bros (Nintendo, 1985) levels [60]. MarioGPT overcomes the issue of data sparsity by using the initial dataset as the starting point for an evolutionary algorithm. Existing levels are selected and then sections of the level are mutated by sampling from the GPT-model and then correcting the border between the re-generated section and the rest of the level with a similarly-trained BERT (i.e. bi-directional) model [61]. This approach produces a large and diverse set of playable levels, despite starting from less than 20 levels.
Both aforementioned GPT-based level generation approaches also show the promise of incorporating natural language instructions to produce conditional level generators, either by prefixing game levels in the training dataset with desired level characteristics [58] or by embedding user instructions and allowing the model to attend to the embedding during generation [60]. It seems likely that more sophisticated techniques such as reinforcement learning with human feedback [2] could produce level generators that are even more capable of accommodating user guidance.
@@ 3.8 Design Assistant
An AI for design assistance can provide several benefits to the creative process. Depending on the type of tool, type of AI, and type of creative process, the AI can minimize development time and cost, reduce human effort, support collaboration among members of a design team or elicit a user’s creativity [62]. So far, in games, most of the AI-powered design assistant tools focus on autocompleting a human’s in-progress design [63] or providing many possible suggestions for the designer to consider [64, 65, 66, 67]. Based on the level of control the AI co-creator [68] has on the process, we identify three levels of assistance:
  • Conceptual assistance where the AI provides high-level guidance which is not game-ready, requiring that the designer adapts and curates the AI output in a way that fits their own vision and the constraints of the game.
  • Procedural assistance where the AI is part of the creative process and through interaction with the user can produce increasingly more final versions of the intended artifact. Here, AI is expected to understand the context of the game for which the content is intended, in order to provide meaningful assistance. However, the AI does not need to produce a final, playable artifact but could instead simply provide the next creative step for discussion with the designer [69]. Moreover, the designer is ultimately responsible for curating and adapting the generated content, as well as deciding when the co-creative process is completed [70].
  • Production assistance where the AI must create the final artifact based on user requirements. This is the closest level to PCG (see Section 3.7), but is different in that the designer remains in control and can refine their specification or reject a created artifact (versus an autonomous generator which directly sends content to the player). As expected, however, the AI operates in a much more constrained space in this scenario as it must account for all other game mechanics (the design of which are presumed finalized) and designer goals which are somehow encoded or presumed via learned designer models [71].
@@
One can argue that existing interfaces with LLMs, Foundation Models and Large Multimodal Models act as design assistants. The designer provides their specifications and receives one (in LLMs) or multiple (in AI image generators) suggestions that they can refine further. Many creatives report using such interfaces for brainstorming and concept development [72], including game developers [73]. However, the applicability of LLMs as design assistants is somewhat limited, reverting only to conceptual assistance. Similarly, their potential for refining an existing idea (i.e. offering procedural assistance) is underexplored, as we discuss in Section 4.
Conceptual assistance is thus the easiest for LLMs, and is the first case explored in games. In [74], design assistance is envisioned in a tool that combines the game description provided by the user with existing knowledge of similar games to suggest possible game features to a designer. The features suggested are fairly generic, few-word guidelines (e.g.: “learn new combat”) which would need extensive design effort and creativity to transform into an implementable and coherent game design.
@@
Since production assistance is also close to traditional PCG pipelines, it is also understandably explored for games. In [75], GPT-3 generates levels from a prompt that describes the level’s features (e.g. width and height) while a human curates and edits the results to ensure playability. This curated set of levels is then used for further fine-tuning, potentially automating the generative process. Instead of generating the final artifact directly, in [76] the LLM is employed as a parser that extrapolates from user requests, expressed in natural language, the high level parameters (level difficulty and size, type of sustainability problem, and simulation goal metrics) required by the PCG system to generate the final game. A collection of candidate game levels are generated from a single user request. Each candidate is then evaluated by an RL agent, and the best game is presented to the user via the Unity game engine.
As noted above, so far research has focused on either LLMs for conceptual assistance (putting significant onus on a human designer) or as production assistance (leveraging a human designer as curator). The conversational nature of LLMs, however, seems particularly well-suited for procedural assistance when designing content; we revisit this missed opportunity in Section 4.

### v2

@@ 3.3 Player Assistant
A somewhat less explored role for LLMs in games is that of a player assistant: an interactive agent intended to enrich or guide the player experience in some way. This could be in the form of a sequence of tutorial-style tips, a character that does not causally interact with the game world at all, or an agent able to interact within the game environment at a similar level as the player. Existing games make use of player assistants in different ways. For example, in The Sims (Electronic Arts, 2000) a disembodied assistant provides tips specific to the game context via dialogue boxes. Civilization VI (Firaxis Games, 2016) uses different assistants with a visual depiction to provide similar advice suggesting the best build option according to their idiosyncratic heuristic; they may thus alleviate some decision-making from the player. In management games, AI may automate menial tasks such as assigning jobs to a planet’s population in Stellaris (Paradox Interactive, 2016); this assistance reduces cognitive load from the player, but the player can always micro-manage this task if they wish.
LLMs are appealing as player assistants given their expressive and conversational capacities. An LLM-based player assistant could plausibly choose an action to suggest to the player, and —more importantly —to form the explanation for this suggestion as a natural language utterance delivered by a disembodied or embodied agent. This utterance could even be accompanied by a corresponding sentiment, and manifested through the assistant’s body stance, gestures and facial expression (in the case of an embodied assistant). The choice of action to suggest to the player could be based on either LLM-powered or heuristic-based methods for finding the best policy or action given the current game context (see Section 3.1). While not intended as a player assistant, this type of exposition is realized by embodied agents in AI people [39] (see Section 3.7). Similarly, LLMs may assist the player by undertaking some minor tasks in the game via a tailored smaller role as “player” within that smaller task description (see Section 3.1). The LLM could extrapolate the policy for such a minor task through conversation with the player, parsing the natural language chat similar to [54]. The potential of LLM-powered player assistants is not explored in current research, to the best of our knowledge. We highlight the potential of this future application in Section 4.
@@ 3.4 Commentator/Reteller
The vision of automated “let’s play”-style commentary generation is not new. It was proposed by Guzdial, Shah and Riedl et al. [56] and implemented via classical machine learning methods, to limited success. Ishigaki et al. [57] trained an LSTM with text, vision and game state input to generate characters for a commentary script in the racing game Assetto Corsa (Kunos Simulazioni, 2013). Results of this approach featured repetitive and context-irrelevant generated text. LSTMs were also used by Li, Gandhi and Harrison [58] to generate text, at a character level, for Getting Over It With Bennett Foddy (Bennett Foddy, 2017), a challenging side-scrolling climbing game.
LLMs for commentary are also explored by Renella and Eger [59], who argue that LLMs could assist game streamers (e.g. on Twitch) while the streamer multitasks gameplay with audience interaction. The authors develop a pipeline for automatically commenting upon League of Legends (Riot Games, 2009) games. They take a multi-phased approach, training a model on hand-annotated data to recognize key events, then prompting ChatGPT to generate zero-shot commentary on these events in the style of a particular (known) fictional character, and finally sending the generated text through the FakeYou [^3] https://fakeyou.com/ API to be voiced in the timbre of this same character. For example, once the event detection model has identified an enemy double kill in a particular frame, ChatGPT responds in the style of Rick Sanchez from Rick and Morty (Cartoon Network, 2013): “What the heck?! That enemy team just got a double kill! I can’t believe it! They must be pretty good! I better watch out for them!” An additional loop buffers detected events —for example, delaying commentary on a double kill in case it should escalate into a triple kill, or prioritizing among a quick barrage of events —and prompts ChatGPT to generate random fillers, such as thanking (fictional) new subscribers.
@@
3.5 Analyst
Another role for LLMs is that of a data analyst. Here, we take this role to primarily analyze player experience and behavior, rather than the broader big-data job title for which some LLMs are naturally suited for [60, 61, 62]. While player modeling is an important aspect of AI research [7] and game development practice [63], LLMs have not received much attention for this purpose. However, the ability of LLMs to make sense of structured data such as code [64, 65] can be a boon for this type of work.
So far, language models (both large and small) have been used for clustering player behaviors. Player2vec [66] trained the Longformer transformer architecture [67], with up to 121 million parameters, on a corpus of game events stored in JSON format which captured player interactions with a casual mobile game. Using dimensionality reduction methods on the latent vectors of the last layer of a pre-trained transformer, the authors discovered eight clusters with player traits that could be useful for market research, such as “lean-in casual economy aware”. In such work, the strength of LLMs is their ability to process structured or unstructured data without the need for extensive data pre-processing. However, as with previous experiments in bottom-up clustering of players [68, 69], the resulting clusters need to be interpreted by expert game analysts (or the game’s designers) to derive some meaningful player types. In this form, current research points to LLMs acting as analyst assistants rather than independent analysts; we revisit future directions in this vein in Section 4.
Representing game logs as text interpreted by an LLM need not be the end-goal for analysis. The LLM representations of game logs can also be used to find common patterns that can be used for other purposes such as gameplay (or gameplay footage) similarity. Indicatively, Trivedi et al. [70] used LLMs on recorded gameplay action logs to establish action similarity. This similarity was in turn applied to train video encodings on gameplay footage of these recorded actions, in order to align the visual latent vectors to the action logs. While the work focused on the computer vision task of representation learning for gameplay footage, better (learned) representations of gameplay pixels would be more generalizable across games in downstream tasks including action recognition or affect modeling; we expand on these downstream tasks in Section 4.
@@ 3.6 Game Master
A Game Master (GM) in tabletop role-playing games (TTRPGs) is the person who creates the plot of a game, its characters, and narrative. GMs wear many hats during the course of the game session [71]; they prepare and adapt stories before sessions, guide gameplay during, and follow up with players afterward [72]. Digital games have mostly pre-scripted stories or level progressions, and their players have a restricted range of affordances, compared to TTRPG players whose actions are only limited by their imagination. Similarly, the story told around the table can take any direction. Since human GMs mostly communicate about the gameworld, story, game state and action resolutions via natural language (although props such as maps, miniatures, hand-outs are also common), the potential of LLMs as a GM is often mentioned both in research circles and TTRPG discussion boards. LLMs as GMs also open the potential for solo play, while a TTRPG requires at least one player and a human GM.
@@
One of the first notable text adventures managed by a fine-tuned version of GPT-2 is AI Dungeon [73]. It is an online [^4] https://play.aidungeon.com/ interactive chat-based storytelling application where the player takes actions through semantic input alone. The LLM continues the story based on the player’s input, in the fashion of a human GM. The game has evolved since its creation to make use of more recent LLM models, which the player can choose from before starting a play session. Different game world settings are also offered, and players are also able to share the stories they create. Similar games have emerged online ever since [^5] https://koboldai.net, https://www.hiddendoor.co, and a freely available code repository, Kobold AI Client [^6] https://github.com/KoboldAI/KoboldAI-Client, allows a local or remote installation of a client for such LLM-run games. Some of these games also use Stable Diffusion text-to-image models [74] to generate visuals accompanying different parts of the narrative. More recent work [75] investigates how different characterizations of the GM and their way of presenting events to the players impacts the overall experience in TTRPGs. This study presented an online interface for a custom game based on a Dungeons & Dragons (TSR, 1974) (D&D) fantasy setting. LLMs were ascribed different GM roles through a number of prompts, and human players participated in 45-minute games (split into 3 equal sub-sessions), with an overall positive response to the automated GM.
@@
In lieu of replacing a human GM, LLMs have also been employed as GM assistants. CALYPSO [76] is a set of tools running on a Discord server which the GM can query either to generate random encounters, brainstorm ideas, or alternatively chat with a fictional character in a D&D TTRPG setting. CALYPSO highlights that hallucinations of GPT-3 can have both positive effects when it generates plausible details not included in descriptions published in the original game manual (e.g. the shapes of creatures’ eyes) and also negative effects when the created details are outright incorrect (e.g. describing the wings of a canonically wingless creature). In addition, the model’s preconditioning to avoid racial bias was found to occasionally prevent it from generating racial details of fantasy creatures in the game. Other work used smaller GPT models to improvise in-game conversations [77] by monitoring and transcribing verbal exchanges between the GM and the players, and attempting to generate appropriate responses. This example was integrated into “Shoelace” [78], which is itself a GM assisting tool helping with content lookup by creating a node-based plan of the game narrative and encounters. The versatility of LLMs given their ability to rapidly process text input paves the way for their integration into the multitude of existing tools and aids for human GMs.
@@ 3.7 Game Mechanic
Games can also be built around a specific mechanic that relies on LLMs, similar to the AI-based game design patterns identified by Treanor et al. [80]. An obvious mechanic revolves around the social interactions facilitated by LLM-powered conversational NPCs. In this vein, the Generative Agents project [43] has employed LLMs to populate a virtual village with 25 characters, enabling them to communicate and engage in social behavior within a sandbox environment. Players were able to interact with these agents via text. The environment state and actions of each agent were stored in a language-based format and summarized in order to retain knowledge for each agent when prompting for its actions. This led to emerging believable social interactions, such as the agents spontaneously inviting other agents to a party which one of them was organizing. Similarly, GoodAI are developing the AI people video game which operates as a sandbox simulation where LLM-powered NPCs “interact with each other and their environment, forming relationships and displaying emotions” [39]. The player can interact with the agents via natural language chat, triggering reactions and potentially disrupting the relationship between NPCs (see Fig. 1).
@@
Another strength of LLMs is language synthesis. Huang at al. [82] used GPT-3 to generate new words from the combination of two user-selected words. These are used to progress in a text-based game scenario, where the player wins by unlocking goal words. Similarly, language synthesis is leveraged by Infinite Craft [^8] https://neal.fun/infinite-craft/, an “alchemy” game, in which the player combines elements to produce new ones (see Fig. 3). In Infinite Craft, the player begins with a set of core elements (water, fire, wind and earth). But while the former have a set of interactions defined manually by the designer, Infinite Craft prompts LLaMA 2 [18] to imagine the product of the combination of these elements [83]. Judging from gameplay, it appears that for each distinct combination, LLaMA is prompted to produce the result only once, with the product stored in a database for future reference. Thus seemingly anything in the language model’s vocabulary might “emerge” from the combination of these elements, including all 50 states [^9] https://x.com/FeralFlex/status/1758332430136615298, “Dream” [^10] https://x.com/slutzsmp/status/1760394123243135169, and the fictional “Super Stonedosaurus Tacosaurus Rex” [^11] https://x.com/pvtspicy/status/1759316982984237139. On occasion, the model can choose to return one of the combined elements, or refuse to combine (e.g. very lengthy or complex) elements.
@@ 3.8 Automated Designer
Prior work in PCG has demonstrated that tile-based game levels can be reliably generated with sequence-based prediction models (e.g. LSTMs) from a modest set of examples, by treating such levels as linear sequences of tile types in raster order [85, 86].
More recently, this approach has been extended to modern LLMs pre-trained on natural language (instead of sequence models trained from the ground up). Todd et al. [87] fine-tuned a GPT-2 model on a large dataset of Sokoban (Thinking Rabbit, 1982) levels and, at test time, samples from the model to produce novel puzzles (see Figure 4). Interestingly, their results indicate that while the GPT-2 model struggles when the fine-tuning dataset is restricted in size, GPT-3 (and, presumably, larger models released since then) are better able to accommodate limited training sets.
@@
Using a similar approach, MarioGPT, trains a GPT-2 model on a relatively small dataset of Super Mario Bros (Nintendo, 1985) levels [89]. MarioGPT overcomes the issue of data sparsity by using the initial dataset as the starting point for an evolutionary algorithm. Existing levels are selected and then sections of the level are mutated by sampling from the GPT model and then correcting the border between the re-generated section and the rest of the level with a similarly-trained BERT (i.e. bi-directional) model [2]. This approach produces a large and diverse set of playable levels, despite starting from less than 20 levels.
Both aforementioned GPT-based level generation approaches also show the promise of incorporating natural language instructions to produce conditional level generators, either by prefixing game levels in the training dataset with desired level characteristics [87] or by embedding user instructions and allowing the model to attend to the embedding during generation [89]. A recent example is Cardistry [90], which used GPT-3.5 to transform a short personal narrative into a set of playing cards. Along with text information, the LLM creates prompts for DALL-E [91], completing the generation pipeline with playing card art.
User requests however are typically fuzzy and can be misinterpreted by the LLM. Hu et al. [54] first “refine” user requests for the generation of Minecraft structures using an LLM, adding knowledge specific to the domain such as block palette and building dimensions. Then, a second LLM is employed to generate the description of the structure that can be interpreted and placed in the game.
Figure 5: A screenshot of a CrawLLM game instance, generated for the theme of Ancient Egypt [92]. Image used with permission.
Unlike the above examples, LLMs can also produce instructions for other LLMs or Foundation Models (FMs) without a user request in the first place. In CrawLLM [92], the Mixtral 8x7B LLM [93] generates the theme, visual style, and even enemy descriptions which act as blueprints to guide additional LLM queries for producing player-facing narrative (e.g. introductory text) or textures and animations via Stable Diffusion [74] for re-theming a dungeon crawler game with card-based combat mechanics. In CrawLLM, the game design (code and card details) are pre-authored by the human, while LLMs and foundation models re-theme the assets to provide a visually and narratively novel experience every time (see Fig. 5).
@@ 3.9 Design Assistant
An AI for design assistance can provide several benefits to the creative process. Depending on the type of tool, type of AI, and type of creative process, the AI can minimize development time and cost, reduce human effort, support collaboration among members of a design , or elicit a user’s creativity [94]. So far, in games, most of the AI-powered design assistant tools focus on autocompleting a human’s in-progress design [95] or providing many possible suggestions for the designer to consider [96, 97, 98, 99, 100, 101]. Ideally, we would want an LLM that can act as a human colleague that we can bounce off ideas to and collaborate with. Such LLMs are still beyond the current state of the art [102]. Existing tools implement co-creating [103] LLMs at different levels of control, which we can explore under the existing Co-Creative Framework for Interaction Design [104]. Focusing on the interaction over the artifact, the LLM can be of conceptual assistance, providing high-level guidance which is not game-ready. This would require that the designer adapts and curates the AI output in a way that fits their own vision and the constraints of the game. This implies that the LLM contributes very little in the actual artifact generation, merely providing new suggestions. By allowing the LLM to also refine or transform the artifact, we see a procedural assistance by the LLM. Interacting with the user, the LLM can produce increasingly more final versions of the intended artifact. The LLM is expected to understand the context of the game for which the content is intended, in order to provide meaningful assistance. However, the LLM does not need to produce a final, playable artifact but could instead simply provide the next creative step for discussion with the designer [105]. Moreover, the designer is ultimately responsible for curating and adapting the generated content, as well as deciding when the co-creative process is completed [106, 107, 108]. Finally, if the LLM is allowed to directly create and alter the artifact based on user requirements, we say it provides production assistance. This is the closest level to PCG (see Section 3.8), but is different in that the designer remains in control and can refine their specification or reject a created artifact (versus an autonomous generator which directly sends content to the player). As expected, however, the AI operates in a much more constrained space in this scenario as it must account for all other game mechanics (the design of which are presumed finalized) and designer goals which are somehow encoded or presumed via learned designer models [109].
@@
One can argue that existing interfaces with LLMs, FMs, and Large Multimodal Models (LMMs) act as design assistants. The designer provides their specifications and receives one (in LLMs) or multiple (in AI image generators) suggestions that they can further refine. Many creatives report using such interfaces for brainstorming and concept development [110], including game developers [111]. However, the applicability of LLMs as design assistants is somewhat limited, reverting only to conceptual assistance. Similarly, their potential for refining an existing idea (i.e. offering procedural assistance) is underexplored, as we discuss in Section 4.
Conceptual assistance is thus the easiest for LLMs, and is the first case explored in games. Charity et al. [112] envision design assistance as a tool that combines the game description provided by the user with existing knowledge of similar games to suggest possible game features back to the user. The features suggested are fairly generic, few-word guidelines (e.g.: “learn new combat”) which would need extensive design effort and creativity to transform into an implementable and coherent game design. When asking an LLM for specific game features to implement in a digital game, however, players would find them less compelling than human-designed ones [113]. The suggestions, however, were still useful for game designers, as they provided a different perspective that could kickstart their creation process. A thorough analysis of strengths and weaknesses of LLMs in this role can be drawn form Project AVA [114], a non-commercial digital game developed at Keywords Studios with the assistance of LLMs and FMs for multiple aspects of the typical game development pipeline. LLMs are shown to help greatly in giving inspiration to the designers (albeit not being creative themselves), provide simple starting code for the game logic, and even provide base assistance to developers with surfacing errors in the code during development. LLMs however often fall short on anything more involved or requiring further domain knowledge, such as requesting code for specific game logic or understanding functionality in UI elements. Similarly, FMs required a lot of tuning by human artists, but provided a solid foundation for concept art and proof-of-concept user interfaces.
@@
Since production assistance is also close to traditional PCG pipelines, it is also understandably explored for games. Nasir and Togelius [115] used GPT-3 to generate levels for the Metavoidal (Yellow Lab Games, 2022) brawler game from a prompt that describes the level’s features (e.g. width and height) while a human curates and edits the results to ensure playability. This curated set of levels is then used for further fine-tuning, potentially automating the generative process. Kelly et al. [116] instead use GPT-4 to generate stories in natural language while abiding to logic constraints, assisting story writers. Instead of generating the final artifact directly, Kumaran et al. [117] use the LLM to parse game level parameters from the user request. The natural language input to the LLM contains information about both what the user is looking for and properties that the level of Future Worlds —a game-based learning exhibit about environmental sustainability [118] —should posses, such as difficulty, size, type of sustainability problem, and simulation goal metrics. A collection of candidate game levels are generated from a single user request. Each candidate is then evaluated by a RL agent, and the best game is presented to the user via the Unity game engine.
Figure 6: A screenshot of the LLMaker digital game content design assistant tool [119]. Image used with permission.
Examples of procedural assistance are very few: LLMaker [119] (see Figure 6) is a chat-only level editor where the user can chat with the LLM and request changes to be made to the level and its content. Natural conversation with the LLM allows the designer to explore the current state of the level (for example, the user may ask “What are the enemies in the entrance room?”) or draw inspiration for the next changes to make (for example, by asking “What kind of enemies should I add to this corridor?”). When the user request a change to be made to the level, the LLM translates their request into a valid function call, ensuring content consistency and adherence to domain constraints [120]. Parameters for the function call that have not been specified by the user are generated by the LLM itself, biased by the existing level and overall user preferences. For example, if the user asks “Change the name of the banshee to Scary Ghost” the LLM will not only change the enemy’s name but also their description to reflect the change, even though this was not specified explicitly by the user (see Fig. 6) In addition the level and its content, LLMaker also employs Stable Diffusion models to generate the graphical assets for the artifacts, based on descriptions generated by the LLM itself. As noted above, so far research has mostly focused on either LLMs for conceptual assistance (putting significant onus on a human designer) or as production assistance (leveraging a human designer as curator). The conversational nature of LLMs, however, seems particularly well-suited for procedural assistance when designing content; we revisit this underexplored area of research in Section 4.